OpenStudy (anonymous):

What are the solutions of 2x2 + 3x= –8?

OpenStudy (anonymous):

quadratic equation

OpenStudy (anonymous):

2x^2+3x+8 = 0 is a^2 +b +c find x use \[(-b \pm \sqrt{b ^{2}-4ac} ) /2a\] \[x _{1,2} = (-3 \pm \sqrt{9-4*2*8} )/ 2*2\] \[x _{1} = (-3 + \sqrt{9-4*2*8} )/ 2*2\] \[= (-3+\sqrt{-57} )/4\] \[x _{2} = (-3 + \sqrt{9-4*2*8} )/ 2*2\] \[= (-3- \sqrt{-57} )/4\]

OpenStudy (anonymous):

no real solution. if u want the imaginaries ^ should b correct